OpenStudy (anonymous):

what is the sum of a, b, c, and d if 3x^3+2x^2+7x+18/x+2 = ax^2+bx+c+d/x+2 ??

OpenStudy (amoodarya):

if you put x=1 you have 3+2+7+18+2=a+b+c+d+2 so a+b+c+d=30 but teh question has logical error

OpenStudy (amoodarya):

are you sure about the type of question ?